The owner of a local motel told you to help him set up his budget with the following information:

Its monthly Fixed Cost amount is $120,000, while its Variable Cost is $15 per room sold.

During the slowest month, he will reduce his ADR to $30 to sell 6,000 rooms. However, the Variable Cost per room will not change. Using the information provided, complete the Pro Forma Income Statements (Operating Budget); and provide your answer to the question below.

    Slowest month

                                                        (6,000 rooms to sell @ lowered ADR of $30)

                                                                                                    Vertical Analysis (%)

           Revenues:                                       $180,000             XXXXXX

           Variable Cost ($15 per room):       __(25)___           __(30)___

           Contribution Margin:                        __(26)___           __(31)___

           Fixed Cost:                                      $120,000             __(32)___

           EBT:                                             __(27)___           XXXXXX

           Tax (40%)                                     __(28)___           XXXXXX

           Net Income:                               __(29)___           __(33)___

1. Provide your answer to the amount of (25)

2. Provide your answer to the amount of (26)

3. Provide your answer to the amount of (27).

4. Provide your answer to the amount of ( 28).

5. Provide your answer to the amount of (29).
Answer #1

1. Variable cost = 6,000*$15 = $90,000

2. Contribution margin = $180,000-90,000 = $90,000

3. EBT = $90,000-120,000 = -$30,000

4. Tax = -$30,000*40% = -$12,000

5. Net income = -$30,000-$12,000) = -$42,000